OpenStudy (secret-ninja):

Solve for x: −3|2x + 6| = −12 x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No solution

OpenStudy (haseeb96):

I 2x+6 I =-12/-3 I 2x+6 I =4 now solve it as 2x +6=4 and 2x +6 =-4 2x=4-6 and 2x =-4-6 2x=-2 and x=-10/2 x=-1 and x=-5 are the correct answer

OpenStudy (haseeb96):

B is the correct answer

OpenStudy (secret-ninja):

Thank you! Can you help me with one more?

OpenStudy (secret-ninja):

Solve for x: -3|x + 7| = -12 x = 5 over 3, x = −19 over 3 x = −3, x = −11 x = −3, x = 11 No solution This one.

OpenStudy (haseeb96):

now write this question on the new ones okay ?